# Beyond Browsing: API-Based Web Agents

### v1

@@ Appendix A Appendix
A.1 WebArena Tasks
WebArena includes the following tasks:
@@ A.1 WebArena Tasks
Gitlab – 180 instances: This website simulates tasks related to project management and version control, where agents perform tasks like opening issues, handling merge requests, or creating repositories. Example query: Submit a merge request for a11yproject.com/redesign branch to be merged into markdown-figure-block branch, assign myself as the reviewer.
@@
Map – 109 instances: For this website, tasks are centered around navigation, trip planning and queries about distances, requiring the agent to retrieve and interpret map-based data, similar to using real-world map services like Google map. Example query: Tell me the full address of all international airports that are within a driving distance of 50 km to Carnegie Mellon University.
@@
Shopping – 187 instances: This dataset represents typical e-commerce tasks, such as searching for products, adding items to carts, and processing transactions. Example query: Change the delivery address for my most recent order to 77 Massachusetts Ave, Cambridge, MA.
@@
Shopping Admin – 182 instances: This setting involves managing backend administrative tasks for an online store, like managing product inventories, processing orders, or viewing sales reports. Example query: Tell me the the number of reviews that our store received by far that mention term "satisfied".
@@
Reddit – 106 instances: Tasks here are similar to interactions with the official Reddit, where agents need to post comments, upvote or down-vote posts, or retrieve information from threads. Example query: Tell me the count of comments that have received more downvotes than upvotes for the user who made the latest post on the Showerthoughts forum.
@@
A.2 API-Based Agent Prompt
A.3 Hybrid Agent Prompt
A.4 Breakdown of Steps and Cost by Website
@@
| Hybrid | 8.1 | 2.0 | 9.4 | 1.7 | 8.2 | 1.3 | 9.0 | 1.4 | 10.5 | 1.0 | 8.0 | 1.9 | 8.9 | 1.5 |

### v2

@@ Appendix A Appendix
A.1 Related Work
The development of AI agents that interact with the web and APIs has garnered significant research attention. Web browsers, serving as the primary interface for interacting with online content, have long been a focus for AI research. Web-based agents that can navigate websites, extract information, and perform tasks autonomously have been studied extensively, especially in the context of LLMs and agents designed to mimic human behavior online.
Web Navigation Agents
Much prior work has centered around agents that perform web-based tasks using browsing actions (Yao et al., 2022; Lai et al., 2024; Koh et al., 2024b; Pan et al., 2024). These agents are particularly effective in environments where human-like interaction with a user interface is necessary (Drouin et al., 2024b). Frameworks such as WebArena have further refined the evaluation of such agents by providing complex and realistic web navigation tasks (Zhou et al., 2024). Our work explores the Hybrid Agent that combines web browsing with API interactions. While prior work primarily focuses on browsing-only agents, we examine how Hybrid Agents can enhance performance by integrating structured API calls with web navigation.
Code Generation Agents and Tool Usage
Another stream of research focuses on agents that interact with online content via application programming interfaces (APIs) (Wang et al., 2024d; Patil et al., 2023; Qin et al., 2023; Yuan et al., 2024; Wang et al., 2024b; Du et al., 2024). In this context, works such as CodeAct have pioneered the development of agents that generate and execute code, including API calls, to perform tasks typically reserved for software engineers (Wang et al., 2024a; Zhang et al., 2024; Tang et al., 2024). These API-Based Agents are optimized for tasks that involve structured data exchanges, allowing them to perform operations more efficiently than traditional web navigation agents (Shen et al., 2024). On the other hand, our work integrates both browsing and API interactions, demonstrating that Hybrid Agents can outperform API-only agents in tasks requiring web navigation. While existing research shows the efficiency of API-Based Agents, our Hybrid Agent dynamically switches between APIs and web browsing to optimize task performance.
A.2 WebArena Tasks
WebArena reproduces the functionality of several commonly-used websites using open-source frameworks, with real-world data imported into the reproduced websites.
WebArena includes tasks related to the following websites:
@@ A.2 WebArena Tasks
Gitlab151515Original Website: https://gitlab.com – 180 instances: This website contains tasks related to project management and version control, where agents perform tasks like opening issues, handling merge requests, or creating repositories. Example query: Submit a merge request for a11yproject.com/redesign branch to be merged into the markdown-figure-block branch, assign myself as the reviewer.
@@
Map161616Original Website: https://www.openstreetmap.org – 109 instances: For this website, tasks are centered around navigation, trip planning and queries about distances, requiring the agent to retrieve and interpret map-based data, similar to using real-world map services like Google map. Example query: Tell me the full address of all international airports that are within a driving distance of 50 km to Carnegie Mellon University.
@@
Shopping171717Developed using Adobe Magento (https://github.com/magento/magento2) – 187 instances: Tasks related to this website represents typical e-commerce tasks, such as searching for products, adding items to carts, and processing transactions. Example query: Change the delivery address for my most recent order to 77 Massachusetts Ave, Cambridge, MA.
@@
Shopping Admin181818Developed using Adobe Magento (https://github.com/magento/magento2) – 182 instances: This setting involves managing backend administrative tasks for an online store, like managing product inventories, processing orders, or viewing sales reports. Example query: Tell me the the number of reviews that our store received by far that mention term “satisfied”.
@@
Reddit191919Deployed Postmill (https://postmill.xyz/), the open-sourced counterpart of Reddit (https://www.reddit.com) – 106 instances: Tasks here are similar to interactions with the official Reddit, where agents need to post comments, upvote or down-vote posts, or retrieve information from threads. Example query: Tell me the count of comments that have received more downvotes than upvotes for the user who made the latest post on the Showerthoughts forum.
@@
A.3 Obtaining APIs of WebArena Websites
Gitlab: we leveraged the open Gitlab REST APIs202020Documentation of all Gitlab APIs could be found at https://docs.gitlab.com/ee/api/rest/., consisting of 988 endpoints. Most of WebArena tasks are covered by these APIs, with only a small fraction of tasks, such as retrieving users’ Gitlab feed token, are not covered by any existing endpoints,
Map: The Map website offers three sets of APIs, each offering distinct functionalities, with a total of 53 endpoints. The first set of APIs, openly available at Nominatim212121The API documentations could be found at https://nominatim.org/release-docs/develop/api/Overview/, offers essential endpoints for geographic searches. The second set of APIs, from Project OSRM222222Documentations of APIs available at https://project-osrm.org/docs/v5.5.1/api, focuses on routing and navigation functionalities. The third set of APIs, available at OpenStreetMap232323API documentations openly available at https://wiki.openstreetmap.org/wiki/API_v0.6, deals primarily with map database operations. This API is rarely used in WebArena tasks but offers capabilities for interacting with OSM data.
Shopping: The e-commerce website uses APIs from the Adobe Commerce API242424APIs documented at https://developer.adobe.com/commerce/webapi/rest/quick-reference/, consisting of 556 endpoints. These endpoints provide support for common shopping tasks such as purchasing products, searching categories, and managing customer accounts.
Shopping Admin: This website shares a common set of APIs with the shopping website. However, this website requires a unique admin token to access the admin-only APIs such as changing the price of products and deleting products from stores.
Reddit: The Reddit tasks in WebArena are based on a self-hosted limited clone of the Reddit website 252525https://codeberg.org/Postmill/Postmill, with limited functionalities as compared to the official site. As a result, all of the available APIs are self-implemented, with a best effort to mimic to official Reddit APIs. This website supports 31 endpoints, which include writing comments and voting posts.
A.4 Additional Analysis
Table 5 documents the percentage of actions of our Hybrid Agent. Across all websites, our Hybrid Agent chooses to do both Browsing and API in the same task at least half of the time.
| Actions | Gitlab | Map | Shopping | Admin | Multi | AVG. | |
| Browsing only | 7.8 | 3.7 | 38.5 | 2.2 | 0 | 8.3 | 12.1 |
| API only | 21.1 | 4.6 | 7.5 | 1.1 | 0 | 10.4 | 7.9 |
| Browsing+API | 71.1 | 91.7 | 54.0 | 96.7 | 82.1 | 81.3 | 80.0 |
| Choices of Action | Gitlab | Map | Shopping | Admin | Multi | AVG. | |
| Browsing only | 7.1(1/14) | 50.0(2/4) | 23.6(17/72) | 50.0(2/4) | 0(0/0) | 25.0(1/4) | 23.5(23/98) |
| API only | 47.4(18/38) | 40.0(2/5) | 21.4(3/14) | 50.0(1/2) | 0.0(0/0) | 20.0(1/5) | 39.1(25/64) |
| Browsing+API | 47.7(61/128) | 46.0(46/100) | 27.7(28/101) | 40.9(72/176) | 51.9(55/106) | 15.4(6/39) | 41.2(268/650) |
Table 6 documents the accuracy of the Hybrid Agent across websites when performing different choices of actions. It shows consistently high accuracy when choosing API only and API+browsing.
@@ A.4 Additional Analysis
| Hybrid | 8.1 | 2.0 | 9.4 | 1.7 | 8.2 | 1.3 | 9.0 | 1.4 | 7.8 | 0.6 | 8.0 | 1.9 | 8.5 | 1.4 |
Steps
Figure 4 demonstrates a scatterplot of the average accuracy of each agent on WebArena over their average steps. The Browsing Agent takes more steps to complete tasks compared to the API-Based Agent on average, while the Hybrid Agent takes the most steps amongst the three agents. This is likely due to the Browsing Agent’s reliance on navigating web interfaces and interacting with visual elements, which involves a sequential and more time consuming processes. The API-Based Agent is the most efficient in terms of steps, as it can directly interact with structured data via APIs, bypassing many of the steps involved in traditional web navigation. The Hybrid Agent, combining both action spaces from the Browsing Agent and the API-Based Agent, takes more steps than both agents.
Costs
Figure 5 demonstrates a scatterplot of the average accuracy of each agent on WebArena over their average costs. The cost of completing tasks shows a different trend. While the Browsing Agent requires more steps, it is much cheaper compared to the API-Based Agent and the Hybrid Agent. This is primarily because the prompts needed for Browsing Agents are much shorter. When browsing, the agent only needs instructions on how to use the web interface and the limited action space around 14 browsing actions. In contrast, API-Based and Hybrid Agents require access to a much larger set of API calls. For example, when interacting with GitLab, the agent is provided with 988 available APIs, leading to much longer prompts and significantly increasing the cost of execution. The cost goes down when the prompt for API calling is shorter. For example, the Reddit website has the least length of API documentation, where its cost is also less than other websites. However, as visualized in Figure 5, the accuracy of the API-Based Agent and the Hybrid Agent is much higher than the Browsing Agent, which makes the increase in cost justifiable due to the significantly improved task performance. The higher cost is offset by the agents’ ability to complete tasks more accurately and efficiently. In the future, this increased cost could potentially be mitigated by methods that retrieve only relevant APIs on the fly.
Error Analysis
We randomly sampled 100 tasks from the WebArena benchmark and performed error analysis on the API-Based Agent. Figure 6 shows the distribution of error categories among these 100 tasks. We found that 33% of the tasks are correctly performed with only API calling, 50% are unsolvable with solely APIs, 6% are incorrect due to incorrect task understanding, and 11% are incorrect due to error in calling APIs such as mal-formatting and wrong input. In other words, among the 50 API solvable tasks, 66% are performed correctly by the API-Based Agent. This showcases the strong capability of the API-Based Agent when given sufficient APIs to solve the task.
Additionally, the average API calls required to solve the API solvable tasks are 2.1 API calls, demonstrating how API calling could reduce operational complexity for web tasks. Although the average number of steps the API-Based Agent took to complete WebArena tasks is 7.8 steps, most of the steps were taken by the agent to retrieve API documentation, resolve errors from it’s previous generations, or verify it’s outputs.
A.5 Case Studies
Case 1
One example where the Hybrid Agent succeeded, while both the API-Based and Browsing Agents failed, involved a task from the Shopping Admin domain. The query was to “delete all negative reviews for Sybil running short”, a product listed in the shopping admin interface. In this instance, the API-Based Agent failed because no relevant API endpoints were available for retrieving or deleting reviews. Similarly, the Browsing Agent failed, as completing this task purely through web navigation required too many steps, as depicted in Figure 7. This complexity made the task challenging for an agent relying solely on web interactions. However, the Hybrid Agent successfully completed the task by leveraging both API and browsing functionalities. An example trace of the Hybrid Agent shown in Figure 7. This case highlights the Hybrid Agent’s ability to efficiently combine API calls with web interactions, allowing it to tackle complex multi-step tasks that would be difficult or impossible for solely browsing or solely API-Based Agents.
Case 2